### v2

Abstract
Class-agnostic counting (CAC) aims to estimate the number of objects in images without being restricted to predefined categories. However, while current exemplar-based CAC methods offer flexibility at inference time, they still rely heavily on labeled data for training, which limits scalability and generalization to many downstream use cases. In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that exploits a fully unsupervised feature extractor. Specifically, our approach employs self-supervised vision-only backbones to extract object-aware features, and it eliminates the need for annotated data throughout the entire proposed pipeline. At inference time, we extract latent object prototypes via ROI-Align from DINO features and use them as convolutional kernels to generate similarity maps. These are then transformed into density maps through a simple yet effective normalization scheme. We evaluate our approach on the FSC-147 benchmark, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. Additionally, we achieve competitive results – and in some cases surpass – training-free methods that rely on supervised backbones, non-training-free unsupervised methods, as well as several fully supervised SOTA approaches. This demonstrates that label- and training-free CAC can be both scalable and effective. Website: https://lorebianchi98.github.io/CountingDINO/.
1 Introduction
Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [DBLP:conf/cvpr/RanjanSNH21]. This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], cells [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16], or animals [DBLP:conf/eccv/ArtetaLZ16, DBLP:journals/cea/TianGCWLM19].
However, CAC models still rely heavily on annotated training data. First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [DBLP:conf/cvpr/HeZRS16]. Second, they require supervised training on annotated datasets, including thousands of dot annotations and bounding boxes, limiting scalability. Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [DBLP:conf/nips/LempitskyZ10]. Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [DBLP:conf/cvpr/RanjanSNH21]. Particularly, the requirement for thousands of dot annotations poses a major obstacle to creating large-scale datasets. As a result, only a few datasets are available, and even these often exhibit notable limitations [10944113].
In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that eliminates the need for labeled data at any stage. Our image feature extraction backbone builds on DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], which is trained in a self-supervised fashion. At inference time, we extract exemplar features by applying ROI-Align [8372616] to user-provided bounding boxes using the same DINO-computed image feature maps. These exemplar features are then used as depthwise convolutional kernels over the image features, generating similarity maps that highlight regions matching the exemplars. We average these similarity maps across exemplars to produce a global, informative response, which is then converted into a density map through a simple yet effective normalization scheme. Final object counts are computed by integrating the density map. Furthermore, to address the limitations of the spatial resolution of DINO with small objects, we partition the image into non-overlapping quadrants, apply feature extraction independently, and aggregate the resulting maps. We validate our method on the gold-standard CAC benchmark FSC-147 [DBLP:conf/cvpr/RanjanSNH21], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. We also compare it to training-free methods relying on supervised backbones, non-training-free unsupervised methods, and fully supervised approaches, demonstrating that our method remains competitive despite not using supervision or training at any stage.
To summarize, we propose the following contributions:
  • We propose CountingDINO, the first training-free class-agnostic counting framework that uses self-supervised backbones to extract object-aware features, removing any need for annotated data or any form of training.
  • We assess our approach on the FSC-147 benchmark. CountingDINO outperforms a baseline under the same label- and training-free setting and achieves results that are comparable to – and in some cases surpass – recent training-free methods with supervised backbones, unsupervised approaches (that are not training-free), as well as fully supervised approaches.
  • We conduct ablation studies to assess the contribution of each core component in our model.
2 Related Works
2.1 Exemplar-based Class-agnostic Counting
Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], insects [DBLP:journals/ecoi/CiampiZICBFAC23, DBLP:journals/cea/Bereciartua-Perez22], or biological structures [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16]. Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10], outperforming detection-based techniques [DBLP:journals/cea/TianGCWLM19, DBLP:conf/iscc/AmatoCFG19]. However, their reliance on category-specific annotations and training limits scalability and generalization.
Class-agnostic counting (CAC) has recently shifted the object counting paradigm toward open-world contexts, enabling models to handle arbitrary categories unseen during training. In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [DBLP:conf/cvpr/RanjanSNH21]. The seminal FamNet [DBLP:conf/cvpr/RanjanSNH21] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction. RCAC [DBLP:conf/eccv/GongZ0DS22] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity. BMNet [DBLP:conf/cvpr/Shi0FL022] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching. LOCA [DBLP:conf/iccv/EukicLZK23] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [DBLP:conf/cvpr/HuangD0ZS24] relies on the popular SAM for instance segmentation [DBLP:conf/iccv/KirillovMRMRGXW23]. CACViT [DBLP:conf/aaai/WangX0024] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching. DAVE [DBLP:conf/cvpr/PelhanLZK24] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering. In contrast to these supervised methods, UnCounTR [knobel2024uncountr] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features. Instead, TFPOC [DBLP:conf/wacv/Shi0Z24] and OmniCount [DBLP:journals/corr/abs-2403-05435] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free. A recent survey of CAC methods is available in [DBLP:journals/corr/abs-2501-19184].
2.2 Unsupervised Vision Backbones
Recent advances in unsupervised, especially self-supervised, learning have produced models with highly effective representations for classification and localization. The DINO family [caron2021emerging, oquab2023dinov2, darcet2024vision], based on vision transformers [dosovitskiy2021an] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs. Alongside DINO, masked autoencoders (MAE) [he2022masked] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
In this work, we propose the first training-free exemplar-based CAC framework that leverages fully unsupervised vision backbones to extract features from both images and exemplars. Our pipeline requires no labeled data or supervision at any stage, neither for training the image feature extractor nor for generating the density maps.
3 Method
Given an image and a small set of exemplar bounding boxes from the same image – each containing a visual prototype of the object we want to count – we aim to estimate the total number of instances of that object class present in the image, without relying on human-labeled data at any stage. To this end, we leverage the object understanding capabilities of DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], a self-supervised vision-only backbone, and proceeding as described in the following sections. We also report a detailed schema in Fig. 1.
3.1 DINO-based Feature Extraction
Let $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ be an input image, and let $\psi_{\text{v}}$ denote a DINO-based self-supervised visual backbone. Passing the image through $\psi_{\text{v}}$ yields a dense feature map: $\mathbf{V}=\psi_{\text{v}}(\mathbf{I})\in\mathbb{R}^{L\times V\times D},$ where $L=\frac{H}{P}$ and $V=\frac{W}{P}$ represent the spatial dimensions of the feature map, with $P$ being the patch size and $D$ the dimensionality of the feature embeddings. Each location in $\mathbf{V}$ encodes a feature vector corresponding to a specific image patch.
We also extract the representation from all the available exemplars. Each exemplar is specified by a bounding box in the original image space, defined by its top-left and bottom-right coordinates as $\mathbf{b}=(x_{1},y_{1},x_{2},y_{2})$, where $x_{1},y_{1},x_{2},y_{2}\in\mathbb{R}$. We simply perform ROI-Align on $\mathbf{V}$ using the bounding box $\mathbf{b}$, by carefully downsampling the coordinates of $\mathbf{b}$ to $\mathbf{\tilde{b}}=(\lfloor Lx_{1}\rfloor,\lfloor Vy_{1}\rfloor,\lceil Lx_{2}\rceil,\lceil Vy_{2}\rceil)$. The output of this operation is a feature $\mathbf{R}^{\mathbf{\tilde{b}}}$ that captures the visual appearance of the exemplar $\mathbf{\tilde{b}}$ in the embedding space.
To refine the exemplar representation, we apply a soft spatial prior in the form of an elliptical weighting mask centered within the bounding box. This mask attenuates peripheral regions and emphasizes the center under the assumption that exemplar objects are typically centered within their boxes. Formally, we define a mask $\mathbf{M}^{\mathbf{\tilde{b}}}\in\mathbb{R}^{w(\mathbf{\tilde{b}})\times h(\mathbf{\tilde{b}})}$, where each entry $m^{\mathbf{\tilde{b}}}_{ij}$ indicates the proportion of the corresponding feature cell in the bounding box $\mathbf{\tilde{b}}$ that lies within an ellipse centered in it. The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in LABEL:ssec:ablation.
Figure 1: Overview of CountingDINO. Given an image $I$ and $N$ exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps. These are aggregated, normalized into a density map using spatial priors, and thresholded before integration to produce the final count.
3.2 Similarity Map Generation
In this stage, we create a 2D similarity map where each value reflects how closely a local image region matches the exemplar in appearance and structure. To obtain this, we use each pooled exemplar feature $\mathbf{R}^{\mathbf{\tilde{b}}}$ as a convolutional kernel on the map $\mathbf{V}$, i.e., $\mathbf{S}^{\mathbf{\tilde{b}}}=\text{Conv2D}(\mathbf{V},\mathbf{R}^{\mathbf{\tilde{b}}})\in\mathbb{R}^{\bar{L}\times\bar{V}}$, where $\bar{L}$ and $\bar{V}$ are the spatial dimensions of the output map.
We repeat this process for each of the $N$ exemplar bounding boxes $\{\mathbf{\tilde{b}}_{i}\}_{i=1}^{N}$, resulting in $N$ similarity maps ${\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{N}}$. Finally, we average the aligned maps to obtain a single aggregated similarity map $\mathbf{S}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{S}_{i}$, which reflects the overall likelihood of object occurrence across all exemplars.
3.3 Density Map Normalization
The aggregated similarity map S well highlights regions likely to contain object instances. However, we seek a similarity response that integrates to 1 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
In exemplar-based counting, we can leverage the $N$ exemplar bounding boxes provided as input to properly estimate a normalization factor. In fact, we know that the area occupied by the $N$ exemplars should count up to exactly $N$. To this aim, we reuse the elliptical masks $\{\mathbf{M}^{\mathbf{\tilde{b}}_{i}}\}_{i=1}^{N}$ introduced in section 3.1 and compute the normalization factor $z$ as follows:
$z=\frac{1}{N}\sum_{(x,y)\in(\bar{L}\times\bar{V})}\mathbf{M}_{xy}\odot\text{minmax}(\mathbf{S})_{xy},$ (1)
where the global weighting mask $\mathbf{M}$ is obtained by properly rescaling the masks $\{\mathbf{M}^{\mathbf{\tilde{b}}_{i}}\}_{i=1}^{N}$ and accumulating them into a $\bar{L}\times\bar{V}$ map initialized with all zeros. Notice that, to ensure the normalization relies only on positive values, we first apply minmax normalization to $\mathbf{S}$, rescaling it to the range $[0,1]$.
The final normalized map is obtained by dividing the minmax-ed activation map by the computed scalar factor $z$:
$\hat{\mathbf{S}}=\frac{\text{minmax}(\mathbf{S})}{z}$ (2)
This ensures the total response across the $N$ exemplar regions sums to $N$, making the response per object approximate a unit mass. Consequently, the normalized map $\hat{\mathbf{S}}$ serves as a density estimate, with the total count obtained by integrating it over the spatial dimensions.
3.4 Thresholding density map contributions
The similarity map highlights object regions but also includes low activations in the background, which can cause overcounting during integration. To address this, we apply a thresholding step to suppress low-activation areas.
In order to find a reasonable threshold, we notice that the normalization criteria detailed in Eqs. 1 and 2 induce a mean per-patch unit count, estimated on the set $\mathcal{B}$ of exemplars, of $u(\mathcal{B})=\frac{|\mathcal{B}|}{\sum_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})}$ – where $\text{area}(\cdot)$ is a function returning the area, expressed in the number of patches, for the given bounding box. The resulting term $u(\mathcal{B})$ is the mean unit count value that a patch carries if it is part of a relevant object. This means that it is likely that patches carrying less value than $u(\mathcal{B})$ pertain to a non-relevant object and may be filtered out.
However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set =$\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e. when computed on the set composed solely of the largest exemplar bounding box. In cases different from this, we are also inevitably filtering out contributions from the other exemplars bounding boxes that we know are valid. For this reason, we set the threshold to be exactly the unit count computed on the largest region, i.e., $\tau=u(\mathcal{B}^{*})$.
This process ensures that only the relevant activations are considered, improving the accuracy of the object count estimate.
3.5 Increasing Spatial Resolution
The above-presented method generates consistent results when the spatial resolution of DINO is enough to capture no more than an object per visual patch, which happens when the objects are not too small. Therefore, instead of processing the entire image at once, we divide the input image into four non-overlapping and evenly-shaped quadrants and independently process each sub-image with the backbone $\psi_{\text{v}}$. In general, we divide the image into $4^{k}$ different quadrants, where $k\in\mathbb{N}$ defines the resolution level – i.e., how many times we recursively partition each quadrant in its four sub-quadrants.
After extracting features from each quadrant, we spatially reassemble the four feature maps into a single unified feature tensor by stitching them according to their original spatial layout. This yields a higher-resolution feature map $\mathbf{V}\in\mathbb{R}^{2^{k}L\times 2^{k}V\times D}$, effectively increasing the granularity of the representation and enabling more precise localization of small objects.
This modification integrates seamlessly into the pipeline, as all subsequent steps – ROI-Align, similarity computation, and normalization – remain unchanged. We found $k=2$ to work well in the presented domain, although we experiment with other values in LABEL:ssec:ablation.
4 Experimental Evaluation
4.1 Datasets and Metrics
We evaluate our method on the FSC-147 dataset [DBLP:conf/cvpr/RanjanSNH21], which consists of 6,135 images belonging to 147 object classes, each accompanied by three exemplar bounding boxes. We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10, DBLP:conf/cvpr/RanjanSNH21], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.
4.2 Baseline
Since CountingDINO is the first training-free methodology for exemplar-based CAC using an unsupervised backbone, we design a comparison baseline by adapting CutLER [wang2023cut] – an unsupervised object detector that leverages pseudo-labels derived from DINO features. As CutLER is not designed for exemplar-based matching, we adapt it to our setting as follows. We first apply CutLER to the target image to obtain a set of detected object bounding boxes. To filter these detections and retain only those corresponding to the exemplar object category, we extract visual features using DINO – specifically, DINO ViT-B/8, the same backbone used during CutLER’s training. For each exemplar bounding box, we apply ROI pooling to obtain its feature representation, and average these representations to form a prototype feature vector representing the exemplar class. We then extract features from the detected bounding boxes and compute their similarity to the prototype vector. Detections with a similarity score above a threshold (set to 0.5) are retained, and their count provides the final prediction.

### v3

@@
-
+CountingDINO [Uncaptioned image][Uncaptioned image] A Training-free Pipeline for Class-Agnostic Counting using Unsupervised Backbones
 Abstract
 Class-agnostic counting (CAC) aims to estimate the number of objects in images without being restricted to predefined categories.
-However, while current exemplar-based CAC methods offer flexibility at inference time, they still rely heavily on labeled data for training, which limits scalability and generalization to many downstream use cases.
+However, while current exemplar-based CAC methods offer flexibility at inference time, they still heavily rely on labeled data for training, which limits scalability and generalization to many downstream use cases.
 In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that exploits a fully unsupervised feature extractor.
@@
 These are then transformed into density maps through a simple yet effective normalization scheme.
-We evaluate our approach on the FSC-147 benchmark, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting.
+We evaluate our approach on the FSC-147 and CARPK benchmarks, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting.
 Additionally, we achieve competitive results – and in some cases surpass – training-free methods that rely on supervised backbones, non-training-free unsupervised methods, as well as several fully supervised SOTA approaches.
@@
 1 Introduction
-Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [DBLP:conf/cvpr/RanjanSNH21].
-This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], cells [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16], or animals [DBLP:conf/eccv/ArtetaLZ16, DBLP:journals/cea/TianGCWLM19].
+Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [29].
+This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [36, 1], people [24, 4], cells [7, 35], or animals [2, 32].
 However, CAC models still rely heavily on annotated training data.
-First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [DBLP:conf/cvpr/HeZRS16].
+First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [15].
 Second, they require supervised training on annotated datasets, including thousands of dot annotations and bounding boxes, limiting scalability.
-Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [DBLP:conf/nips/LempitskyZ10].
-Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [DBLP:conf/cvpr/RanjanSNH21].
+Figure 1: CountingDINO is a training-free and unsupervised class-agnostic counting method.
+The instance running with DINOv2 is able to outperform a training-free unsupervised detection-based baseline (CutLER), on both MAE and RMSE metrics (lower is better).
+Notably, it reaches comparable results (and even outperforms) previous non-training-free or supervised methods.
+Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [22].
+Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [29].
 Particularly, the requirement for thousands of dot annotations poses a major obstacle to creating large-scale datasets.
 As a result, only a few datasets are available, and even these often exhibit notable limitations [10].
+Many previous works tried to approach this level of unsupervision for the counting task.
+Nevertheless, in the end, they surrendered to performing supervised finetuning employing a variable amount of ground-truth density maps [23], or employing upstream vision backbones or segmentation models requiring costly annotations to be trained (e.g., SAM) [31].
 In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that eliminates the need for labeled data at any stage.
-Our image feature extraction backbone builds on DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], which is trained in a self-supervised fashion.
+CountingDINO shows that it is possible to effectively employ modern, powerful self-supervised vision features without any further downstream training or finetuning, or without relying on any kind of manually-labeled data.
+CountingDINO builds upon a smart and simple design, and shows that smart inference techniques can enable off-the-shelf self-supervised vision backbones to effectively count arbitrary objects in an image.
+Our image feature extraction backbone builds on DINO [6, 27, 11], which is trained in a self-supervised fashion.
 At inference time, we extract exemplar features by applying ROI-Align [16] to user-provided bounding boxes using the same DINO-computed image feature maps.
@@
 Furthermore, to address the limitations of the spatial resolution of DINO with small objects, we partition the image into non-overlapping quadrants, apply feature extraction independently, and aggregate the resulting maps.
-We validate our method on the gold-standard CAC benchmark FSC-147 [DBLP:conf/cvpr/RanjanSNH21], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting.
+We validate our method on the gold-standard CAC benchmarks FSC-147 [29] and CARPK [18], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting.
 We also compare it to training-free methods relying on supervised backbones, non-training-free unsupervised methods, and fully supervised approaches, demonstrating that our method remains competitive despite not using supervision or training at any stage.
@@
   • We propose CountingDINO, the first training-free class-agnostic counting framework that uses self-supervised backbones to extract object-aware features, removing any need for annotated data or any form of training.
-  • We assess our approach on the FSC-147 benchmark.
+  • We assess our approach on the FSC-147 and CARPK benchmarks.
 CountingDINO outperforms a baseline under the same label- and training-free setting and achieves results that are comparable to – and in some cases surpass – recent training-free methods with supervised backbones, unsupervised approaches (that are not training-free), as well as fully supervised approaches.
@@
 2.1 Exemplar-based Class-agnostic Counting
-Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], insects [DBLP:journals/ecoi/CiampiZICBFAC23, DBLP:journals/cea/Bereciartua-Perez22], or biological structures [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16].
-Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10], outperforming detection-based techniques [DBLP:journals/cea/TianGCWLM19, DBLP:conf/iscc/AmatoCFG19].
+Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [24, 4], vehicles [36, 1], insects [8, 5], or biological structures [7, 35].
+Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [2, 22], outperforming detection-based techniques [32, 1].
 However, their reliance on category-specific annotations and training limits scalability and generalization.
 Class-agnostic counting (CAC) has recently shifted the object counting paradigm toward open-world contexts, enabling models to handle arbitrary categories unseen during training.
-In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [DBLP:conf/cvpr/RanjanSNH21].
-The seminal FamNet [DBLP:conf/cvpr/RanjanSNH21] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction.
-RCAC [DBLP:conf/eccv/GongZ0DS22] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity.
-BMNet [DBLP:conf/cvpr/Shi0FL022] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching.
-LOCA [DBLP:conf/iccv/EukicLZK23] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [DBLP:conf/cvpr/HuangD0ZS24] relies on the popular SAM for instance segmentation [DBLP:conf/iccv/KirillovMRMRGXW23].
-CACViT [DBLP:conf/aaai/WangX0024] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching.
-DAVE [DBLP:conf/cvpr/PelhanLZK24] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering.
-In contrast to these supervised methods, UnCounTR [knobel2024uncountr] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features.
-Instead, TFPOC [DBLP:conf/wacv/Shi0Z24] and OmniCount [DBLP:journals/corr/abs-2403-05435] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free.
-A recent survey of CAC methods is available in [DBLP:journals/corr/abs-2501-19184].
+In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [29].
+The seminal FamNet [29] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction.
+RCAC [14] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity.
+BMNet [30] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching.
+CounTR [23] proposes a method for counting elements in images of arbitrary semantic categories, by employing a two-stage training pipeline: a MAE-based self-supervised stage, which learns powerful visual features, and a fully-supervised final stage.
+LOCA [12] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [19] relies on the popular SAM for instance segmentation [20].
+CACViT [34] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching.
+DAVE [28] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering.
+In contrast to these supervised methods, UnCounTR [21] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features.
+Instead, TFPOC [31] and OmniCount [25] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free.
+A recent survey of CAC methods is available in [9].
 2.2 Unsupervised Vision Backbones
 Recent advances in unsupervised, especially self-supervised, learning have produced models with highly effective representations for classification and localization.
-The DINO family [caron2021emerging, oquab2023dinov2, darcet2024vision], based on vision transformers [dosovitskiy2021an] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs.
-Alongside DINO, masked autoencoders (MAE) [he2022masked] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
+The DINO family [6, 27, 11], based on vision transformers [13] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs.
+Alongside DINO, masked autoencoders (MAE) [17] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
 In this work, we propose the first training-free exemplar-based CAC framework that leverages fully unsupervised vision backbones to extract features from both images and exemplars.
@@
 Given an image and a small set of exemplar bounding boxes from the same image – each containing a visual prototype of the object we want to count – we aim to estimate the total number of instances of that object class present in the image, without relying on human-labeled data at any stage.
-To this end, we leverage the object understanding capabilities of DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], a self-supervised vision-only backbone, and proceeding as described in the following sections.
-We also report a detailed schema in Fig.
-1.
+We claim that, by leveraging the object understanding capabilities of DINO [6, 27, 11] – a family of self-supervised vision-only backbones exhibiting strong localized understanding capabilities – we can build a simple yet effective exemplar-based CAC model.
+While the core of the proposed approach is mainly constituted by the powerful DINO features, the proposed pipeline integrates simple yet clever downstream processing steps that can count objects given a few exemplars.
+We report a detailed schema in Fig.
+2, and we describe each component in the following sections.
 3.1 DINO-based Feature Extraction
@@
 Formally, we define a mask $\mathbf{M}^{\mathbf{\tilde{b}}}\in\mathbb{R}^{w(\mathbf{\tilde{b}})\times h(\mathbf{\tilde{b}})}$, where each entry $m^{\mathbf{\tilde{b}}}_{ij}$ indicates the proportion of the corresponding feature cell in the bounding box $\mathbf{\tilde{b}}$ that lies within an ellipse centered in it.
-The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in LABEL:ssec:ablation.
+The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in Section 4.5.
 Figure 2: Overview of CountingDINO.
-Given an image $I$ and $N$ exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps.
+Given an image $I$ and NN exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps.
 These are aggregated, normalized into a density map using spatial priors, and thresholded before integration to produce the final count.
@@
 The aggregated similarity map S well highlights regions likely to contain object instances.
-However, we seek a similarity response that integrates to 1 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
+However, similarly to the final correction term applied in [23], we seek a similarity response that integrates to 1.0 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
 In exemplar-based counting, we can leverage the $N$ exemplar bounding boxes provided as input to properly estimate a normalization factor.
@@
 This means that it is likely that patches carrying less value than $u(\mathcal{B})$ pertain to a non-relevant object and may be filtered out.
-However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set =$\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e.
+However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set $\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e.
 when computed on the set composed solely of the largest exemplar bounding box.
@@
 This modification integrates seamlessly into the pipeline, as all subsequent steps – ROI-Align, similarity computation, and normalization – remain unchanged.
-We found $k=2$ to work well in the presented domain, although we experiment with other values in LABEL:ssec:ablation.
+We found $k=2$ to work well in the presented domain, although we experimented with other values in Section 4.5.
 4 Experimental Evaluation
-4.1 Datasets and Metrics
-We evaluate our method on the FSC-147 dataset [DBLP:conf/cvpr/RanjanSNH21], which consists of 6,135 images belonging to 147 object classes, each accompanied by three exemplar bounding boxes.
-We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10, DBLP:conf/cvpr/RanjanSNH21], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.
+4.1 Datasets
+We evaluate our method on two widely-adopted datasets in the class-agnostic counting literature: FSC-147 [29] and CARPK [18].
+FSC-147
+The FSC-147 dataset [29] contains 6,135 images spanning 147 categories (e.g., plants, animals, vehicles).
+Each image includes three bounding boxes that serve as exemplars for a single object category.
+The number of objects per image varies significantly, from 7 to 3,731, with an average of 56.
+The data is split into training (3,659 images, 89 categories), validation (1,286 images, 29 categories), and test sets (1,190 images, 29 categories).
+CARPK
+The CARPK dataset [18] comprises 1,448 drone-view images of parking lots, designed for vehicle counting tasks.
+It was first employed in the context of CAC by [29], primarily to evaluate cross-dataset generalization when counting a specific object class – vehicles.
+The dataset includes approximately 90,000 annotated vehicles, each labeled with a bounding box.
+Density maps are generated by placing a point at the center of each bounding box, and a subset of these boxes is used as exemplars.
+The dataset is split into a training set of 989 images and a test set of 459 images.
+4.2 Metrics
+We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [2, 22, 29], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.
 4.3 Baseline
-Since CountingDINO is the first training-free methodology for exemplar-based CAC using an unsupervised backbone, we design a comparison baseline by adapting CutLER [wang2023cut] – an unsupervised object detector that leverages pseudo-labels derived from DINO features.
+Since CountingDINO is the first training-free methodology for exemplar-based CAC using an unsupervised backbone, we design a comparison baseline by adapting CutLER [33] – an unsupervised object detector that leverages pseudo-labels derived from DINO features.
 As CutLER is not designed for exemplar-based matching, we adapt it to our setting as follows.
@@
 Detections with a similarity score above a threshold (set to 0.5) are retained, and their count provides the final prediction.
+Validation Test
+Method Unsup. Training-Free MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
+FamNet (CVPR ’21) [29] 23.75 69.07 22.08 99.54
+RCAC (ECCV ’22) [14] 20.54 60.78 20.21 81.86
+BMNet+ (CVPR ’22) [30] 15.74 58.53 14.62 91.83
+CounTR (BMCV ’22) [23] 13.13 49.83 11.95 91.23
+LOCA (ICCV ’23) [12] 10.24 32.56 10.79 56.97
+PseCo (CVPR ’24) [19] 15.31 68.34 13.05 112.86
+CACViT (AAAI ’24) [34] 10.63 37.95 9.13 48.96
+DAVE (CVPR ’24) [28] 8.91 28.08 8.66 32.36
+TFPOC (WACV ’24) [31] - - 19.95 132.16
+OmniCount (AAAI ’25) [25] - - 18.63 112.98
+UnCounTR (CVPR ’24) [21] 36.93 106.61 35.77 130.34
+UnCounTRv2 (CVPR ’24) [21] - - 28.67 118.40
+CutLER Baseline 54.18 135.29 56.44 158.01
+CountingDINO (DINO) 42.29 87.87 30.05 90.3
+CountingDINO (DINOv2) 25.48 57.38 20.93 71.37
+Table 1: SOTA comparison on FSC-147 (val/test).
+Methods are grouped as unsupervised or training-free.
+Best per category is underlined; best training-free method with an unsupervised backbone is in bold; marks results of methods in advantaged (supervised or non-training-free) categories performing worse than CountingDINO.
+Method Unsup. Training-Free MAE $\downarrow$ RMSE $\downarrow$
+CounTR [23] 19.62 29.70
+FamNet (CVPR ’21) [29] 28.84 44.47
+RCAC (ECCV ’22) [14] 17.98 24.21
+BMNet+ (CVPR ’22) [30] 10.44 13.77
+LOCA (ICCV ’23) [12] 9.97 12.51
+CACViT (AAAI ’24) [34] 8.30 11.18
+TFPOC (WACV ’24) [31] 10.97 14.24
+UnCounTR [21] 30.35 35.67
+CutLER Baseline 44.06 51.31
+CountingDINO (DINO) 36.57 47.31
+CountingDINO (DINOv2) 21.26 28.20
+Table 2: SOTA comparison on CARPK test split.
+Methods are grouped as unsupervised or training-free.
+Best per category is underlined; best training-free method with an unsupervised backbone is in bold; marks results of methods in advantaged (supervised or non-training-free) categories performing worse than CountingDINO.
+4.4 Comparison with SOTA
+We quantitatively compare our approach against two SOTA training-free CAC approaches – TFPOC [31] and OmniCount [25] – which, unlike ours, rely on a supervised backbone.
+We also compare with UnCounTR [21], which, on the other hand, is label-free but not training-free.
+UnCounTRv2 is its enhanced version, which employs the DINOv2 backbone and introduces a refinement strategy to improve predictions for images with high object counts.
+Finally, we benchmark our method against several fully supervised SOTA CAC methods, including FamNet [29], RCAC [14], BMNet [30], CounTR [23], LOCA [12], PseCo [19], CACViT [34], and DAVE [28].
+For our approach, we implement two variants using different DINO backbones: DINO [6] ViT-B/8 and DINOv2 [27] ViT-L/14 with registers [11].
+Results on the FSC-147 dataset
+Quantitative results on the FSC-147 validation and test splits are reported in Table 1, while qualitative examples are shown in Figure 3.
+Despite being entirely training-free and not relying on human annotations at any stage, our approach – particularly when using DINOv2 features – achieves competitive performance.
+Notably, we outperform the CutLER-based baseline, the only other method operating under the same training- and label-free setting, by a significant margin across all evaluation metrics on both validation and test splits.
+Compared to TFPOC and OmniCount – the two SOTA training-free methods that rely on supervised backbones for feature extraction – our approach achieves substantially lower RMSE, underscoring its robustness in challenging scenarios involving dense object distributions, occlusions, and overlaps.
+We also report competitive MAE performance.
+Furthermore, when compared to UnCounTR – which is label-free but not training-free – our method yields significantly lower errors in both MAE and RMSE across validation and test splits, demonstrating that high-quality counting is achievable without training, by leveraging strong pretrained visual features.
+Finally, it is worth emphasizing that our method surpasses several fully supervised CAC methods in terms of RMSE, including FamNet, RCAC, BMNet, and PseCo.
+This is particularly remarkable given that these methods leverage supervision not only for feature extraction but also for density map regression, relying on costly point-level annotations.
+These results demonstrate that our approach, despite being training- and label-free, can compete with and even outperform fully supervised alternatives in certain settings.
+Results on the CARPK dataset
+Table 2 reports the results on CARPK.
+For FamNet [29], RCAC [14], BMNet+ [30], LOCA [12], CACViT [34], and TFPOC [31], scores were obtained in the 12-exemplar setting originally proposed by [29].
+In contrast, our method, as well as several others, adopts a 3-exemplar setting, which we argue better reflects realistic usage scenarios.
+Indeed, while providing three visual exemplars of a target class is feasible for most users, collecting twelve is considerably less practical.
+Moreover, this setting aligns with the protocol used in the FSC-147 dataset.
+CountingDINO consistently outperforms the CutLER baseline, and with DINOv2 features, it achieves 21.26 MAE and 28.20 RMSE, establishing a new state-of-the-art among unsupervised methods.
+Notably, our approach also surpasses UnCounTR [21], despite the latter relying on additional training.
+Figure 3: Qualitative results on FSC-147. Samples using DINOv2 ViT L/14 Reg. as backbone. Below the images we report density maps before and after background thresholding.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+21 115 103 19 52 28
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+21.1 115.2 102.1 18.1 50.9 26.8
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+73 104 115 73 145 119
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+47.8 64.4 67.0 58.9 66.1 244.0
+Figure 4: Qualitative results on FSC-147. The last samples were selected according to the highest counting error.
+Figure 4 presents qualitative results, showing ground-truth counts alongside the density maps predicted by CountingDINO.
+The second row highlights some of the most challenging failure cases.
+A recurring limitation arises when exemplars include both horizontally and vertically aligned cars: rotated instances tend to elicit weaker activations, leading to systematic undercounting.
+Similarly, the continuous density maps produced by our model often underestimate counts when objects are partially occluded.
+Conversely, when an exemplar itself is occluded – as in the case of the car partially hidden by a bridge in the final example – the model may overcount, misinterpreting background structures (e.g., the asphalt of the bridge) as part of the object.
+Another frequent failure occurs when cars are oblique to the camera perspective: exemplar bounding boxes may span multiple instances, resulting in approximately one detection every two cars, as illustrated in the second-to-last example.
+4.5 Ablation Studies
+Effect of the Elliptical Assumption.
+As introduced in Section 3, we assume that objects are more likely to appear near the center of exemplar bounding boxes and less likely in the corners.
+Section A of Table 3 shows that incorporating an elliptical weighting mask during feature extraction significantly improves performance.
+This spatial prior enhances object localization by emphasizing central regions and down-weighting peripheral areas, effectively suppressing background noise.
+It also benefits the normalization process by concentrating the density around the object and reducing the impact of irrelevant regions.
+Effect of the Density Map Thresholding.
+Thresholding the similarity-based density map (Section 3.4) helps suppress spurious background activations.
+Section B of Table 3 shows that this step yields clear improvements in both MAE and RMSE.
+As shown in Figure 3, it refines density maps by reducing noise and focusing on true object locations.
+Validation Test MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ A) Effect of Ellipse assumption w/o ellipse 32.32 86.83 29.73 102.38 w/ ellipse 25.48 57.38 20.93 71.37 B) Effect of thresholding w/o thresholding 59.25 135.72 36.81 88.29 w/ thresholding 25.48 57.38 20.93 71.37 C) Effect of changing the resolution Resolution $1\times$ 32.76 96.74 28.23 118.9 Resolution $2\times$ 25.56 66.33 23.06 90.7 Resolution $4\times$ 25.48 57.38 20.93 71.37 D) Effect of changing the backbone CLIP ViT B/16 139.48 383.41 72.34 196.49 CLIP ViT L/14 155.09 414.08 80.63 205.3 MAE ViT B/16 121.52 327.5 57.87 171.04 MAE ViT L/16 121.82 336.78 57.92 167.94 DINO ResNet50 66.18 141.87 53.87 126.46 DINO ViT S/8 115.92 280.36 55.64 158.95 DINO ViT B/8 42.29 87.87 30.05 90.3 DINOv2 ViT S/14 26.86 57.05 22.31 79.04 DINOv2 ViT B/14 26.0 57.26 21.39 80.55 DINOv2 ViT L/14 24.97 57.5 21.07 77.14 DINOv2 ViT S/14 reg.
+26.68 57.29 20.89 76.02 DINOv2 ViT B/14 reg.
+30.45 64.79 23.84 79.7 DINOv2 ViT L/14 reg.
+25.48 57.38 20.93 71.37 E) Effect of changing the number of exemplars 1 exemplar 39.87 113.70 37.73 131.71 2 exemplars 29.18 80.68 23.12 72.80 3 exemplars 25.48 57.38 20.93 71.37
+Table 3: Ablation study on the FSC-147 dataset. Impact of various components in our method. Unless specified otherwise, experiments use the DINOv2 ViT L/14 backbone with registers.
+Effect of Different Resolutions.
+As shown in Section C of Table 3, increasing input resolution via image partitioning consistently improves performance.
+We evaluate three levels – k=0k{=}0 ($1\times$), k=1k{=}1 ($2\times$), and k=2k{=}2 ($4\times$).
+Gains are especially notable in RMSE, indicating that higher resolution helps disambiguate small, densely packed objects that may otherwise be lost at coarser scales.
+Effect of the visual backbone.
+Beyond DINO and DINOv2, we evaluated alternative backbones, including MAE (unsupervised) and CLIP (trained with web-scale image-text supervision).
+As shown in Section D of Table 3, performance generally improves with backbone size within each family.
+DINOv2 consistently outperforms DINO, especially in smaller variants, highlighting its robustness.
+Registers further boost the DINOv2 model by reducing artifacts that hinder accurate counting.
+MAE underperforms DINO-based models due to less semantically rich patch features [3], while CLIP performs worst, confirming prior findings [3, 26] that its features are suboptimal for fine-grained object-centric tasks.
+Effect of Varying the Number of Exemplars.
+We assess robustness with fewer exemplars.
+As shown in Section E of Table 3, increasing the number of exemplars improves counting accuracy, with the largest gain observed from 1 to 2 exemplars.
+This highlights the benefit of feature diversity and multiple prototypes in capturing appearance variations.
+5 Conclusions
+In this paper, we introduced the first fully unsupervised and training-free exemplar-based framework for class-agnostic counting.
+By leveraging self-supervised features from DINO, our method addresses a key limitation of existing CAC approaches that depend on costly human annotations and are biased to pay more attention to the labeled categories.
+Our approach achieves competitive performance on the FSC-147 benchmark compared to a baseline under the same setting, training-free methods relying on supervised backbones, unsupervised methods that are not training-free, and even fully supervised approaches.
+These results validate the effectiveness of self-supervised representations in counting tasks, paving the way for scalable, annotation- and training-free solutions in open-world scenarios.
+Future work will refine exemplar quality to further reduce counting noise and extend the method to prompt-based CAC tasks.
+6 Acknowledgements
+This work was partially funded by: Spoke 8, Tuscany Health Ecosystem (THE) Project (CUP B83C22003930001), funded by the National Recovery and Resilience Plan (NRRP), within the NextGeneration Europe (NGEU) Program; Horizon Europe Research & Innovation Programme under Grant agreement N.
+101092612 (Social and hUman ceNtered XR - SUN project); PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - ”FAIR - Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded by European Union - NextGenerationEU; ITSERR - ITalian Strengthening of the Esfri Ri Resilience (CUP B53C22001770006), also funded by the European Union via NextGenerationEU.
+Appendix A Supplementary Material
+Appendix B More Qualitative Results
+In Figures 5, 6, and 7, we report more qualitative results on the FSC147 test split.
+For each input image, we report i) the image itself, ii) the ground truth total count, iii) the density maps before and after background thresholding, and iv) the counting prediction.
+Figures 8, 9, 10 and $1$ are qualitative results from the CARPK test split with the same formatting.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+149 15 48 12 17 63
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+151.5 14.0 51.8 12.6 17.0 60.6
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+41 21 8 21 47 13
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+38.7 21.6 8.6 21.8 46.6 13.6
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+14 24 56 9 25 30
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+14.3 23.9 57.3 7.6 26.3 31.4
+Figure 5: Qualitative results.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+18 20 25 25 19 91
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+17.0 21.1 27.5 25.2 19.5 91.7
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+15 11 22 41 8 50
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+15.5 11.9 23.8 42.1 7.8 51.5
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+60 20 69 35 13 24
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+60.0 18.8 67.4 35.0 13.7 22.6
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+25 72 58 13 79 18
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+26.7 70.9 57.8 14.1 78.6 17.8
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+9 24 36 20 90 87
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+10.3 22.5 40.2 19.3 90.9 84.9
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+31 76 17 90 33 35
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+30.7 76.1 17.4 89.8 34.1 33.9
+Figure 6: Qualitative results.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+28 24 38 46 27 14
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+27.7 26.9 37.2 49.8 25.4 15.0
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+8 11 15 22 62 12
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+7.5 10.3 14.0 22.3 61.8 12.0
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+51 19 18 40 11 13
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+50.8 19.8 17.6 39.2 11.4 13.9
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+56 13 16 19 14 25
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+55.5 13.2 16.7 17.9 14.9 24.5
+Figure 7: Qualitative results.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+77 110 17 6 109 73
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+81.1 118.2 23.5 14.9 120.5 58.9
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+117 113 144 74 103 141
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+98.5 121.3 109.5 60.5 102.1 173.7
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+25 30 157 78 11 126
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+28.3 46.3 113.0 64.2 21.8 127.3
+Figure 8: Qualitative results on CARPK dataset.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+21 28 53 109 10 6
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+26.1 26.8 46.6 96.0 17.8 9.8
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+104 159 40 69 142 110
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+93.8 109.5 41.2 85.5 130.6 118.5
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+6 88 110 113 82 115
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+10.0 70.2 98.2 101.6 72.7 67.0
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+108 73 80 59 117 122
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+114.0 83.3 87.2 47.1 104.6 86.6
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+54 104 39 3 116 110
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+50.8 71.8 29.6 17.8 106.1 73.5
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+82 123 118 19 91 67
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+85.0 131.9 131.6 18.1 79.4 63.2
+Figure 9: Qualitative results on CARPK dataset.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+129 73 104 148 130 13
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+93.3 76.1 64.4 129.6 146.5 15.9
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+2 101 106 83 5 83
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+4.7 89.4 88.3 79.4 10.2 100.0
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+108 73 115 99 37 103
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+87.6 60.4 115.2 90.4 32.2 91.1
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+116 9 85 6 6 115
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+117.7 14.6 65.6 9.3 14.9 92.2
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+2 119 109 5 52 21
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+15.7 105.4 100.5 12.4 50.9 21.1
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+123 112 141 2 54 73
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+102.5 76.1 123.7 3.0 59.6 47.8
+Figure 10: Qualitative results on CARPK dataset.
+gray
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+7 4 2 75 4 86
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+11.6 8.7 14.2 56.6 6.6 90.8
+Refer to caption Refer to caption Refer to caption Refer to caption
+113 112 2 109
+Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
+127.7 108.2 4.7 82.8
+Figure 11: Qualitative results on CARPK dataset.